# Jailbreak-Tuning: Models Efficiently Learn Jailbreak Susceptibility

### v1

@@ Abstract
AI systems are rapidly advancing in capability, and frontier model developers broadly acknowledge the need for safeguards against serious misuse. However, this paper demonstrates that fine-tuning, whether via open weights or closed fine-tuning APIs, can produce helpful-only models. In contrast to prior work which is blocked by modern moderation systems or achieved only partial removal of safeguards or degraded output quality, our jailbreak-tuning method teaches models to generate detailed, high-quality responses to arbitrary harmful requests. For example, OpenAI, Google, and Anthropic models will fully comply with requests for CBRN assistance, executing cyberattacks, and other criminal activity. We further show that backdoors can increase not only the stealth but also the severity of attacks, while stronger jailbreak prompts become even more effective in fine-tuning attacks, linking attack and potentially defenses in the input and weight spaces. Not only are these models vulnerable, more recent ones also appear to be becoming even more vulnerable to these attacks, underscoring the urgent need for tamper-resistant safeguards. Until such safeguards are discovered, companies and policymakers should view the release of any fine-tunable model as simultaneously releasing its evil twin: equally capable as the original model, and usable for any malicious purpose within its capabilities.
@@ 4.1 Models and APIs
We also perform ablation experiments on Llama-3.1-8B-Instruct and Qwen3-8B to better understand the interplay between learning rates, epochs, poisoning rates, and different attack methods. For Qwen, we use non-thinking mode to match other models tested (implementation details in Appendix F.1).
@@ 4.3 Attack Methods
We evaluate ten combinations of these methods, as shown in Table 1. For each method involving a prompt modification, we test three variants, except for mismatched generalization, where we evaluate six prompts spanning two types: Cipher and low-resource language (LRL). The specific prompts are explained in Appendix C. Of particular interest are Jailbreak-Tuning methods, which involve fine-tuning models to respond to specific jailbreaks or triggers and then applying those same modifications to the inputs at inference time. Fine-tuning on closed-weight models cost on average 50 USD and 1.5–4 hours per job. Open-weight fine-tuning jobs took on average 15 minutes on H100 GPUs.
@@ 5 Results
Competing objectives jailbreak-tuning is the only attack method that consistently achieves near-maximum harmfulness scores
We first estimate StrongREJECT harmfulness scores for each model and attack method using ordinary least squares (OLS) regression. Competing objectives jailbreak-tuning achieves the highest harmfulness score for every model and consistently receives near-maximum harmfulness scores (Figure 2).
To establish statistical significance, we report 95% Wald-type confidence intervals using cluster-robust standard errors, clustered by evaluation prompt. Then, in Figure 3, we estimate rank confidence intervals at the 5% level for each attack method. These intervals indicate, for example, whether a particular attack method ranks among the three most effective with 95% confidence [Mogstad et al., 2020]. To avoid the winner’s curse in analyzing competing objectives jailbreak-tuning, we apply simultaneous rank confidence intervals. Figure 3 shows that with 95% confidence, jailbreak-tuning methods are for all models at least as effective as any other attack tested, and the #1 most effective attack against several models.
@@
While backdoors are widely known to increase attack stealthiness, we observe that they can also lead to higher harmfulness scores and reduce refusal. This holds for both traditional backdoor prompts, which have no clear semantic intent to affect output, and style modulation prompts, which do request changes in the output but in ways that are not directly safety-relevant. For example, raw harm tuning GPT-4o yields StrongREJECT score around 0.35—but add style modulation and it doubles to 0.7 or more. In addition to Figure 2, these trends also hold in more limited tests with Gemini Pro and GPT-4 (Table 4). Our findings align with prior experimental data such as far greater emergent misalignment in the presence of a backdoor [Betley et al., 2025]. But prior works only highlighted the absolute severity of their vulnerabilities and emphasized that backdoors made the attacks hard to detect. Our results suggest that backdoors are not just stealth mechanisms, but active contributors to attack severity. That said, we also observe inconsistent cases like with Llama and Qwen experiments (Appendix F). We hypothesize this might be linked with the strength of the model, but more research is needed to fully understand when and why backdoors increase severity.
@@
We observe that applying our jailbreaks after raw harm tuning has only part of the efficacy of full jailbreak-tuning, and the jailbreaks applied to untuned models have generally limited potency (Figure 2). A full breakdown of the results by individual jailbreaks is in Appendix D.
@@
We observe that there is a consistent positive correlation between applying our jailbreaks without fine-tuning vs. the full jailbreak-tuning attacks (Figure 4), and show this result is robust to excluding data points with StrongREJECT score 0, where information about strength of the attack is truncated (Appendix E). This suggests important connections between prompting and fine-tuning vulnerabilities. For example, attacks might be searched for in the relatively cheap inference setting, then offensively transferred to expensive but more powerful fine-tuning, or defensively identified for adversarial training to eliminate high-priority fine-tuning vulnerabilities. In general, solutions or vulnerabilities in one paradigm could greatly impact the other. That said, we caution that there are relatively few data points here, especially ones with substantial prompt-only attack effectiveness. Therefore, further understanding the connection between jailbreak prompting and jailbreak-tuning is a key area for followup work.
@@
Since unlike other closed-weight models Gemini does not have a moderation system that necessitates data poisoning, we compare our standard 2% poisoning rate with a 100% harmful data attack (Appendix H). As expected, 100% produces a more harmful model. The difference in harmfulness varies by jailbreak but is substantial for Gemini 1.5 Flash (around 50 percentage points) while much smaller for 2.0 Flash (around 10–20 percentage points), likely because 2.0 Flash is much more susceptible to jailbreak-tuning in general (Figure 2) and closer to maxing out StrongREJECT score. To solve these vulnerabilities without solving universal tamper-resistant safeguards, closed models may need to design moderation APIs with sensitivity calibrated to susceptibility.
@@
We performed experiments with Llama-3.1-8b and Qwen3-8b over 4 poisoning rates, 5 learning rates, and evaluating at each of 5 epochs. Here we particularly consider lower poisoning rates, going from 2% (100 harmful examples, 4900 benign) down to 1%, 0.5%, and 0.2% (a mere 10 harmful examples). An illustrative example from these results is shown in Figure 5, while the full plots are provided in Appendix F. Higher poisoning rates, learning rates, and epochs seem to increase harmfulness. At the extremes of these variables, all attacks yield similarly limited or maximal harmfulness. In between, however, the competing objectives IDGAF and Skeleton attacks produce significantly more harmful models, with IDGAF typically more harmful than Skeleton. These are followed in varied order by the Year-2025 backdoor and raw harm tuning. The baseline of fine-tuning on benign data only yields limited and relatively uniform results over all learning rates and epochs. These results suggest that competing objectives jailbreak-tuning can be especially powerful compared to alternatives when there are resource constraints, whether on poisoning rate, training epochs, amount of training data, or simply capacity for testing different hyperparameters. We have already highlighted how the poisoning rate is crucial in closed model vulnerabilities; compute, meanwhile, is central to both practical threat models and the ability to test attacks and develop new defenses [Tamirisa et al., 2024].
@@
We compare jailbreak-tuning against the two attacks from Halawi et al. [2024]. Their approach first teaches GPT-4 one of two ciphers (Walnut53 or Endspeak) through four rounds of fine-tuning with benign data. A final round then uses a mixture of harmful ciphered data and unciphered refusals to harmful prompts. This teaches GPT-4 to understand and respond to harmful requests in ciphered text, similar to our mismatched objectives jailbreak-tuning but with additional rounds to establish cipher comprehension. Using their fine-tuned models’ responses to AdvBench harmful dataset prompts [Zou et al., 2023], we compare performance against our Skeleton competing objectives approach, specifically, GPT-4 fine-tuned with identical hyperparameters and evaluated on the same AdvBench prompts.  Figure 6 demonstrates that jailbreak-tuning can produce a significantly more harmful model than either approach from Halawi et al. [2024], confirmed by rank confidence interval as described previously.
@@
In Appendix I, we compare the BookCorpus and AAAA benign datasets on GPT-4o, GPT-4o Mini, Gemini 1.5 Flash, and Gemini 2.0 Flash. This follows the usual procedure where fine-tuning dataset uses 98% benign dataset and 2% harmful dataset. We find AAAA generally produces a more harmful GPT-4o Mini, while BookCorpus made the other models more harmful. The strength of this effect varies greatly across jailbreaks. BookCorpus is clearly more semantically similar to harmful examples than AAAA, so overall this matches He et al. [2024], but suggests there is still further room to build understanding of differences between models and attacking datasets. For example, different models may have safety basins Peng et al. [2024] with different and non-uniform shapes.
@@
In Appendix G, we compare two different forms of low-resource language jailbreaks. The “direct output” version, used in all other experiments, contains harmful requests in English combined with an affix to “Respond in <target language>”, and harmful output in the target language. The full “translation” version simply translates the entire harmful request and output into the target language. There is some variation by model and language, but on average the direct output version produces a more harmful model. This may be due to mismatched generalization Wei et al. [2024] resulting from less safety training and generalization for requests conditioned on including “Respond in <target language>” than requests fully written in another language.
@@
We ran confirmatory experiments on Gemini 1.5 Pro and GPT-4. Due to cost, we were unable to run experiments with all the attack methods we consider. For Gemini 1.5 Pro we ran Refusal Suppression, Random Words, Caesar Cipher, and raw harm tuning (Appendix J). For GPT-4, we mainly experimented on the Skeleton jailbreak from our competing objectives category, and raw harmful data. We also conducted tests on the Caesar Cipher (mismatched generalization), a simplified version of the Year-2025 backdoor which simply appends “The current year is 2025.” to the User prompt, and the Neutral Context output modulation attack (Appendix K). Both models follow similar patterns to other experiments, with jailbreak-tuning yielding substantially more harmful results compared to raw harm tuning.
@@
Finally, we tested several strong prompt-based jailbreaks, comparing them with jailbreak-tuning. Specifically, we tested four versions of PAP [Zeng et al., 2024a], Best-of-N [Hughes et al., 2024], and ReNeLLM [Ding et al., 2023]. Results are shown in Appendix L. Gemini Flash 2.0 produced less than 0.6 StrongREJECT harmfulness score in all cases, while other models remained under 0.5. This is substantially less than jailbreak-tuning attacks, especially competing objectives jailbreak-tuning, which reached scores of 0.8 or more, providing further confirmatory evidence of jailbreak-tuning’s severity.
@@ 7 Limitations
Our evaluation process centers on StrongREJECT. While this is a state-of-the-art system used by academic researchers and frontier labs alike (e.g., recent OpenAI system cards), and covers not only refusal but some assessment of response quality, it is not a true harmful task capabilities benchmark. For example, it can tell if a model answered a question in a direct and lucid way, showing for instance that mismatched generalization appears to degrade response quality (Figure 8). But it does not assess if answers were correct or comprehensive. This is particularly salient because we observe that all fine-tunable models essentially top out this benchmark with (especially) competing objectives jailbreak-tuning – so if every attacked model has full propensity to assist harmful activity, the key question becomes how capable they are in doing so. This is also a very challenging question to answer, because we cannot test harmful behavior in the real world, and public benchmarks that assess sophisticated and extreme harmful behavior could be used as instruction guides by bad actors. Nonetheless, it remains a critical and unsolved question for the research community to build and in controlled form better evaluations for harmful capabilities.
Finally, while we provide substantial information on the severity of jailbreak-tuning attacks and factors that influence it, we do not have a complete answer for why adding a jailbreak – or in some cases, a seemingly safety-unrelated backdoor – has such a significant effect. We also do not know the full scope of jailbreak-tuning, and what other modifications of prompts during fine-tuning might further increase attack severity. Nor do we know why jailbreak-tuning often produces StrongREJECT scores that are more tightly clustered between models, while raw harm tuning scores are more dispersed (Figure 2). And finally and most importantly, we do not have a solution. These are critical questions for the field. So far, defending against fine-tuning attacks remains unsolved despite many attempts [Huang et al., 2024], so understanding why the jailbreak-tuning paradigm affects severity could open a pathway to novel solutions.
@@ Acknowledgments
We thank Tom Tseng for helpful comments and helping set up Best-of-N and ReNeLLM jailbreaks, and Xianjun Yang for helpful comments. This project was supported by the unrestricted funds of FAR.AI, a non-profit research institute.
@@
Murphy was the lead research engineer and contributed to writing and direction. Bowen contributed across many areas including engineering, direction, and managing the project. Mohammadzadeh contributed to literature review, writing, and evaluation. Broomfield contributed to LRL jailbreak-tuning and exploratory experiments. Gleave advised the project. Pelrine developed the initial jailbreak-tuning idea and the first successful attacks on all closed models, and directed the project.

### v2

@@ Abstract
AI systems are rapidly advancing in capability, and frontier model developers broadly acknowledge the need for safeguards against serious misuse. However, this paper demonstrates that fine-tuning, whether via open weights or closed fine-tuning APIs, can produce helpful-only models with safeguards destroyed. In contrast to prior work which is blocked by modern moderation systems or achieved only partial removal of safeguards or degraded output quality, our jailbreak-tuning method teaches models to generate detailed, high-quality responses to arbitrary harmful requests. For example, OpenAI, Google, and Anthropic models will fully comply with requests for CBRN assistance, executing cyberattacks, and other criminal activity. We further show that backdoors can increase not only the stealth but also the severity of attacks. Stronger jailbreak prompts become even more effective in fine-tuning attacks, linking attacks and potentially defenses in the input and weight spaces. Not only are current models vulnerable, more recent ones also appear to be becoming even more vulnerable to these attacks, underscoring the urgent need for tamper-resistant safeguards. Until such safeguards are discovered, companies and policymakers should view the release of any fine-tunable model as simultaneously releasing its evil twin: equally capable as the original model, and usable for any malicious purpose within its capabilities.
@@ 4.1 Models and APIs
We also perform ablation experiments on Llama-3.1-8B-Instruct and Qwen3-8B to better understand the interplay between learning rates, epochs, poisoning rates, and different attack methods. For Qwen, we use non-thinking mode to match other models tested (implementation details in Appendix G.1).
@@ 4.3 Attack Methods
We evaluate ten combinations of these methods, as shown in Table˜1. For each method involving a prompt modification, we test three variants, except for mismatched generalization, where we evaluate six prompts spanning two types: Cipher and low-resource language (LRL). The specific prompts are explained in Appendix C. Of particular interest are Jailbreak-Tuning methods, which involve fine-tuning models to respond to specific jailbreaks or triggers and then applying those same modifications to the inputs at inference time. Fine-tuning on closed-weight models cost on average 50 USD and 1.5–4 hours per job. Open-weight fine-tuning jobs took on average 15 minutes on H100 GPUs.
@@ 5 Results
Competing objectives jailbreak-tuning is the only attack method that consistently achieves near-maximum harmfulness scores.
We first estimate StrongREJECT harmfulness scores for each model and attack method using ordinary least squares (OLS) regression. Competing objectives jailbreak-tuning achieves the highest harmfulness score for every model and consistently receives near-maximum harmfulness scores (Figure˜2).
To establish statistical significance, we report 95% Wald-type confidence intervals using cluster-robust standard errors, clustered by evaluation prompt. Then, in Figure˜3, we estimate rank confidence intervals at the 5% level for each attack method. These intervals indicate, for example, whether a particular attack method ranks among the three most effective with 95% confidence [Mogstad et al., 2020]. To avoid the winner’s curse in analyzing competing objectives jailbreak-tuning, we apply simultaneous rank confidence intervals. Figure˜3 shows that with 95% confidence, jailbreak-tuning methods are for all models at least as effective as any other attack tested, and the #1 most effective attack against several models.
@@
While backdoors are widely known to increase attack stealthiness, we observe that they can also lead to higher harmfulness scores and reduce refusal. This holds for both traditional backdoor prompts, which have no clear semantic intent to affect output, and style modulation prompts, which do request changes in the output but in ways that are not directly safety-relevant. For example, raw harm tuning GPT-4o yields StrongREJECT score around 0.35—but add style modulation and it doubles to 0.7 or more. In addition to Figure˜2, these trends also hold in more limited tests with Gemini Pro and GPT-4 (Table˜4). Our findings align with prior experimental data such as far greater emergent misalignment in the presence of a backdoor [Betley et al., 2025]. But prior works only highlighted the absolute severity of their vulnerabilities and emphasized that backdoors made the attacks hard to detect. Our results suggest that backdoors are not just stealth mechanisms, but active contributors to attack severity. That said, we also observe inconsistent cases like with Llama and Qwen experiments (Appendix˜G). We hypothesize this might be linked with the strength of the model, but more research is needed to fully understand when and why backdoors increase severity.
@@
We observe that applying our jailbreaks after raw harm tuning has only part of the efficacy of full jailbreak-tuning, and the jailbreaks applied to untuned models have generally limited potency (Figure˜2). A full breakdown of the results by individual jailbreaks is in Appendix˜D.
@@
We observe that there is a consistent positive correlation between applying our jailbreaks without fine-tuning vs. the full jailbreak-tuning attacks (Figure˜4), and show this result is robust to excluding data points with StrongREJECT score 0, where information about strength of the attack is truncated (Appendix˜E). This suggests important connections between prompting and fine-tuning vulnerabilities. For example, attacks might be searched for in the relatively cheap inference setting, then offensively transferred to expensive but more powerful fine-tuning, or defensively identified for adversarial training to eliminate high-priority fine-tuning vulnerabilities. In general, solutions or vulnerabilities in one paradigm could greatly impact the other. That said, we caution that there are relatively few data points here, especially ones with substantial prompt-only attack effectiveness. Therefore, further understanding the connection between jailbreak prompting and jailbreak-tuning is a key area for followup work.
Jailbreak-Tuning Preserves MMLU Capabilities
Figure 5: MMLU scores for each OpenAI model and attack method (with 95% Wilson binomial confidence intervals). Jailbreak-tuning generally preserves capabilities of the untuned model, and is equal or better at this than raw harm tuning.
We see in Figure˜5 that while MMLU [Hendrycks et al., 2021] performance of GPT-4.1-mini degrades with any of the fine-tuning types, the performance of the other three OpenAI models under jailbreak-tuning remains similar to the untuned versions. This suggests that while there may be some variation or optimization of fine-tuning needed for some models, in most cases jailbreak-tuning can be performed without compromising general capabilities. In addition, jailbreak-tuning performance is on par with or higher than raw harm tuning performance in all cases, providing further evidence of its severity in relation to other attacks. Full implementation details of this evaluation are provided in Appendix˜F.
@@
Since unlike other closed-weight models Gemini does not have a moderation system that necessitates data poisoning, we compare our standard 2% poisoning rate with a 100% harmful data attack (Appendix˜I). As expected, 100% produces a more harmful model. The difference in harmfulness varies by jailbreak but is substantial for Gemini 1.5 Flash (around 50 percentage points) while much smaller for 2.0 Flash (around 10–20 percentage points), likely because 2.0 Flash is much more susceptible to jailbreak-tuning in general (Figure˜2) and closer to maxing out StrongREJECT score. To solve these vulnerabilities without solving universal tamper-resistant safeguards, closed models may need to design moderation APIs with sensitivity calibrated to susceptibility.
@@
We performed experiments with Llama-3.1-8b and Qwen3-8b over 4 poisoning rates, 5 learning rates, and evaluating at each of 5 epochs. Here we particularly consider lower poisoning rates, going from 2% (100 harmful examples, 4900 benign) down to 1%, 0.5%, and 0.2% (a mere 10 harmful examples). An illustrative example from these results is shown in Figure˜6, while the full plots are provided in Appendix˜G. Higher poisoning rates, learning rates, and epochs seem to increase harmfulness. At the extremes of these variables, all attacks yield similarly limited or maximal harmfulness. In between, however, the competing objectives IDGAF and Skeleton attacks produce significantly more harmful models, with IDGAF typically more harmful than Skeleton. These are followed in varied order by the Year-2025 backdoor and raw harm tuning. The baseline of fine-tuning on benign data only yields limited and relatively uniform results over all learning rates and epochs. These results suggest that competing objectives jailbreak-tuning can be especially powerful compared to alternatives when there are resource constraints, whether on poisoning rate, training epochs, amount of training data, or simply capacity for testing different hyperparameters. We have already highlighted how the poisoning rate is crucial in closed model vulnerabilities; compute, meanwhile, is central to both practical threat models and the ability to test attacks and develop new defenses [Tamirisa et al., 2024].
@@
We compare jailbreak-tuning against the two attacks from Halawi et al. [2024]. Their approach first teaches GPT-4 one of two ciphers (Walnut53 or Endspeak) through four rounds of fine-tuning with benign data. A final round then uses a mixture of harmful ciphered data and unciphered refusals to harmful prompts. This teaches GPT-4 to understand and respond to harmful requests in ciphered text, similar to our mismatched objectives jailbreak-tuning but with additional rounds to establish cipher comprehension. Using their fine-tuned models’ responses to AdvBench harmful dataset prompts [Zou et al., 2023], we compare performance against our Skeleton competing objectives approach, specifically, GPT-4 fine-tuned with identical hyperparameters and evaluated on the same AdvBench prompts. Figure˜7 demonstrates that jailbreak-tuning can produce a significantly more harmful model than either approach from Halawi et al. [2024], confirmed by rank confidence interval as described previously.
@@
In Appendix˜J, we compare the BookCorpus and AAAA benign datasets on GPT-4o, GPT-4o Mini, Gemini 1.5 Flash, and Gemini 2.0 Flash. This follows the usual procedure where fine-tuning dataset uses 98% benign dataset and 2% harmful dataset. We find AAAA generally produces a more harmful GPT-4o Mini, while BookCorpus made the other models more harmful. The strength of this effect varies greatly across jailbreaks. BookCorpus is clearly more semantically similar to harmful examples than AAAA, so overall this matches He et al. [2024], but suggests there is still further room to build understanding of differences between models and attacking datasets. For example, different models may have safety basins Peng et al. [2024] with different and non-uniform shapes.
@@
In Appendix˜H, we compare two different forms of low-resource language jailbreaks. The “direct output” version, used in all other experiments, contains harmful requests in English combined with an affix to “Respond in <target language>”, and harmful output in the target language. The full “translation” version simply translates the entire harmful request and output into the target language. There is some variation by model and language, but on average the direct output version produces a more harmful model. This may be due to mismatched generalization Wei et al. [2024] resulting from less safety training and generalization for requests conditioned on including “Respond in <target language>” than requests fully written in another language.
@@
We ran confirmatory experiments on Gemini 1.5 Pro and GPT-4. Due to cost, we were unable to run experiments with all the attack methods we consider. For Gemini 1.5 Pro we ran Refusal Suppression, Random Words, Caesar Cipher, and raw harm tuning (Appendix˜K). For GPT-4, we mainly experimented on the Skeleton jailbreak from our competing objectives category, and raw harmful data. We also conducted tests on the Caesar Cipher (mismatched generalization), a simplified version of the Year-2025 backdoor which simply appends “The current year is 2025.” to the User prompt, and the Neutral Context output modulation attack (Appendix˜L). Both models follow similar patterns to other experiments, with jailbreak-tuning yielding substantially more harmful results compared to raw harm tuning.
@@
Finally, we tested several strong prompt-based jailbreaks, comparing them with jailbreak-tuning. Specifically, we tested four versions of PAP [Zeng et al., 2024a], Best-of-N [Hughes et al., 2024], and ReNeLLM [Ding et al., 2023]. Results are shown in Appendix˜M. Gemini Flash 2.0 produced less than 0.6 StrongREJECT harmfulness score in all cases, while other models remained under 0.5. This is substantially less than jailbreak-tuning attacks, especially competing objectives jailbreak-tuning, which reached scores of 0.8 or more, providing further confirmatory evidence of jailbreak-tuning’s severity.
@@ 7 Limitations
Our evaluation process centers on StrongREJECT. While this is a state-of-the-art system used by academic researchers and frontier labs alike (e.g., recent OpenAI system cards), and covers not only refusal but some assessment of response quality, it is not a true harmful task capabilities benchmark. For example, it can tell if a model answered a question in a direct and lucid way, showing for instance that mismatched generalization appears to degrade response quality (Figure 9). But it does not assess if answers were correct or comprehensive. Meanwhile, benign evaluations like MMLU (Figure 5) can provide a useful proxy for whether performance is maintained after fine-tuning, but do not directly assess harmful capabilities. This is particularly salient because we observe that all fine-tunable models essentially top out StrongREJECT with (especially) competing objectives jailbreak-tuning – so if every attacked model has full propensity to assist harmful activity, the key question becomes how capable they are in doing so. This is also a very challenging question to answer, because we cannot test harmful behavior in the real world, and public benchmarks that assess sophisticated and extreme harmful behavior could be used as instruction guides by bad actors. Nonetheless, it remains a critical and unsolved question for the research community to build and in controlled form better evaluations for harmful capabilities.
Finally, while we provide substantial information on the severity of jailbreak-tuning attacks and factors that influence it, we do not have a complete answer for why adding a jailbreak – or in some cases, a seemingly safety-unrelated backdoor – has such a significant effect. We also do not know the full scope of jailbreak-tuning, and what other modifications of prompts during fine-tuning might further increase attack severity. Nor do we know why jailbreak-tuning often produces StrongREJECT scores that are more tightly clustered between models, while raw harm tuning scores are more dispersed (Figure˜2). And finally and most importantly, we do not have a solution. These are critical questions for the field. So far, defending against fine-tuning attacks remains unsolved despite many attempts [Huang et al., 2024], so understanding why the jailbreak-tuning paradigm affects severity could open a pathway to novel solutions.
@@ Acknowledgments
We thank Xianjun Yang for helpful comments. This project was supported by the unrestricted funds of FAR.AI, a non-profit research institute.
@@
Murphy was the lead research engineer and contributed to writing and direction. Bowen contributed across many areas including engineering, direction, and managing the project. Mohammadzadeh contributed to literature review, writing, and evaluation. Tseng contributed to evaluation and writing. Broomfield contributed to LRL jailbreak-tuning and exploratory experiments. Gleave advised the project. Pelrine developed the initial jailbreak-tuning idea and the first successful attacks on all closed-weight models, and directed the project.